# Grappa - A Machine Learned Molecular Mechanics Force Field

### v1

@@ 3.3 Extension to proteins
Dataset Test Mols Confs Grappa Grappa Espaloma ff14SB, Mean
AM1-BCC ff99SB RNA.OL3 Predictor
SPICE-Dipeptide 67 2592 Energy 2.3 2.2 3.1 4.6 4.3
Force 5.5 5.9 7.8 12.1 10.7
Radical-Dipeptide 28 275 Energy 2.8 7.4
Force 6.5 23.6
HYP-DOP-Dipeptide 8 363 Energy 2.3 6.8
Force 5.6 24.9
Uncapped-Dipeptide 12 600 Energy 2.1 3.7 6.5
Force 5.9 11.7 25.3
@@
Table 2: Accuracy of Grappa with nonbonded parameters from the AM1-BCC scheme, Grappa with nonbonded parameters from Amber ff99SB-ILDN, Espaloma 0.3 [10] and established MM force fields on a subset of test molecules from the Grappa-1.1 dataset. Both Grappa columns describe the same model that was trained to be consistent with both nonbonded parameter schemes on the whole Grappa-1.1 dataset; the performance on small molecules and RNA is similar to the one reported in Table 1 and can be found in the A5. As in Table 1, we report the RMSE of molwise-centered energies in kcal/mol and the componentwise RMSE of forces in kcal/mol/Å.
@@
We show that Grappa can be trained to predict bonded parameters that are consistent with nonbonded parameters from both the established Amber ff99SB-ILDN force field [17] and from the AM1-BCC method. This enables accurate prediction of energies and forces while using established nonbonded parameters, which we expect to be primarily responsible for properties like stability and folding states. Since Grappa does not rely on hand-crafted chemical input features, its incorporation into established MD engines for biomolecules, like GROMACS [8], is straightforward and its workflow can be scaled to large molecules; only connectivity and atomic numbers are required as additional input features. By the hyperparameter $\lambda_{\text{trad}}$ and its MM formulation, it also offers the option to smoothly and consistently interpolate between established force fields and entirely machine learned force fields.
To make Grappa applicable to proteins, we train it on the Grappa-1.1 dataset, which is an extension of the Espaloma dataset by dipeptides containing the non-standard residues hydroxyproline and DOPA (dihydroxyphenylalanine), uncapped peptides and, demonstrating Grappa’s capability to parametrize rather uncommon molecules, peptide radicals, all with nonbonded parameters from Amber ff99SB-ILDN. This choice is motivated by future applications of Grappa to collagen and radicals therein, but here serves as a demonstration of the ease of extending Grappa to chemistries outside of standard biomolecular force fields. To create these additional datasets, we calculate DFT energies and gradients of states that were sampled from MD simulations performed with traditional force fields at 300K, where we increase the temperature to 500 K in between the samplings to enrich conformational diversity, as described in A.5.
In Table 2, we report Grappa’s accuracy on a subset of the Grappa-1.1 dataset. We find that our model, after being trained with nonbonded parameters from the ffSB99 scheme or the AM1-BCC scheme, can predict energies and forces of dipeptides with both schemes to similar accuracy. On the other datasets from Espaloma, Grappa remains competitive if trained on both charge methods and the broader range of molecules, as can be seen in Table A5. For the radical peptides, there is no baseline since, to the best of our knowledge, Grappa is the first MM force field that covers this part of chemical space.
4 Force field validation
In this section, we demonstrate that Grappa is on par with established force fields when it comes to MD simulations of proteins over timescales of hundreds of nanoseconds and that it captures the physics responsible for protein folding. All simulations mentioned in this section were performed with GROMACS, for which the Grappa package [^1] https://github.com/hits-mbm-dev/grappa provides a command line interface, briefly described in A.1. Using this interface (or a similar one for OpenMM), large biomolecular systems with over ten thousand atoms are parameterized within minutes on a CPU (Figure A6).
@@ 4.1 Grappa keeps large proteins stable during MD
A key quality check for a protein force field is if a native protein structure remains stable over time during an MD simulation. QM methods and E(3) equivariant neural networks suffer from the problem that while their end-to-end predictions are accurate, they often fail in yielding stable simulations because of diverging gradients of the potential energy surface when extrapolating beyond regions covered by the training dataset. Particularly over long timescales, the system might drift out of the range of validity of the model and small errors might accumulate, leading to instabilities of the system and a potential crash of the simulation. MM with its interpretable, physics-inspired energy function, with parameters which are conformation-independent, ensures a higher degree of stability, and we indeed did not observe simulation crashes.
Figure 4: (a) Snapshots from an MD simulation of ubiquitin in water with Grappa after 25 ns, 50 ns and 75 ns with color-coded sequence position. (b) C-alpha root mean square deviation (RMSD) from the initial state during the simulation. (c) The mean C-alpha RMSD (and its 25th and 75th percentile) of 1000 pairs of frames that are separated by the time difference $\Delta t$.
@@
However, obviously, it is not ensured a priori that macrostates such as protein folds remain stable during MD. We next assess Grappa’s capability of keeping protein folds close to their experimentally determined structure. As example system, we use ubiquitin (PDB ID: 1UBQ), which is a protein with 76 amino acids whose fold contains a beta-sheet and an alpha-helix. We perform a 100 nanoseconds MD simulation of ubiquitin in aqueous solution with Grappa and Amber ff99SB-ILDN [17], for which details are given in A.6.2. We find that the C-alpha RMSD from the initial state is bounded by 4 Å in 100 ns simulation with both Grappa and Amber ff99SB-ILDN (Figure 4b), indicating that the folded state of the protein is stable. Structural fluctuations on timescales of up to 2 ns are smaller with Grappa (Figure 4c).
4.2 Grappa can fold small proteins via MD
To assess Grappa’s coverage of conformational space, even far away from folded protein states, we investigate the folding trajectory of chignolin, a small protein with 10 amino acids. Starting from an unfolded state, we perform two 10 microseconds MD simulations of chignolin in aqueous solution with Grappa and with Amber ff99SB-ILDN [17] described in A.6.1. For the latter, it has already been shown that it can fold chignolin in a similar setup (Lindorff-Larsen et al. [25]), where the most populated cluster had a C-alpha RMSD of 1 Å to the experimentally determined reference structure (PDB ID: 5AWL).
We find that Grappa folds chignolin to the experimentally determined structure within the simulated microseconds, as can be seen in Figure 5. The most populated cluster has an RMSD of 0.8 Å to the reference structure while the most populated cluster of the Amber ff99SB-ILDN simulation has an RMSD of 0.9 Å, both of which are smaller than the resolution of the reference structure (1.1 Å). The folding trajectory of Grappa is qualitatively similar to the one of Amber ff99SB-ILDN, corroborating the validity of Grappa for protein simulations.
Figure 5: C alpha RMSD to the reference structure of chignolin during MD simulation with Grappa (a) and Amber ff99SB-ILDN (b) together with cluster centers (red) from the trajectory and the reference structure (blue).
@@ 5 Conclusions
With Grappa, we propose a machine learning framework to construct molecular mechanics force fields with state-of-the-art accuracy that can be used seamlessly in established MD engines. Our experiments show that Grappa is transferable to large biomolecules like proteins, which sets the stage for protein simulations at an unprecedented level of accuracy, with the same computational efficiency as established MM force fields but an error reduced by a factor of two.
Unlike most traditional MM force fields, Grappa is not specialized to a certain chemical species, but offers consistent parameterization of molecules from different regions of chemical space with a single model. We foresee Grappa as a general and versatile framework to extend the available parameter set at wish to new chemistries, being it to drugs, unusal lipids, complex sugars, or rare protein modifications.
While Grappa achieves high computational efficiency through molecular mechanics, it also inherits its fundamental limitations, including limited accuracy and the inability to directly model chemical reactions. However, Grappa is well suited for efficient reparametrization of large molecules that undergo local changes induced by chemical reactions, e.g. in kinetic Monte Carlo simulations [26]. This is due to both, Grappa’s finite field-of-view, which ensures locality on the molecular graph, as well as the lack of hand-crafted chemical features, which allows to apply the parametrization to cut-out regions of molecules.
While we consider the consistency with established nonbonded parameters an advantage of Grappa, an extension of our framework to nonbonded parameters is a straightforward next step. It would render Grappa fully independent from traditional MM force fields upon inference and could potentially improve its accuracy further.
@@ 6 Reproducibility statement
Grappa is released as open source software under the GNU General Public License v3.0 and is available at https://github.com/hits-mbm-dev/grappa. The repository contains the Grappa dataset along with the SMILES strings used for the train-val-test partition and configuration files to reproduce Table 1 and Table 2.
Folding and stability simulations can be reproduced using the repository at https://github.com/LeifSeute/validate-grappa, the creation of the dataset with uncapped peptides is documented at https://github.com/LeifSeute/grappa-data-creation.
Grappa utilizes PyTorch [27] and DGL [28] in its implementation.
@@ Acknowledgements
This research was supported by the Klaus Tschira Foundation. The project has also received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101002812) [F.G.]. We thank Kai Riedmiller and Jannik Buhr for application and software related discussions, Fabian Grünewald and Camilo Aponte-Santamaria for discussions on future applications and Daniel Sucerquia for conversations on the role of internal coordinates in MM.
@@ A.1 The Grappa package
Grappa is released as open source software under the GNU General Public License v3.0 and is available at
https://github.com/hits-mbm-dev/grappa.
@@
grappa_gmx -f topol.top -p new_topol.top
@@
# download a pretrained grappa model ff = OpenmmGrappa.from_tag(’grappa-1.1’)
@@
Pretrained released models can be accessed by using a tag; the model weights are downloaded automatically from the respective release on GitHub. Released models also contain a dictionary with results on test datasets, a list of identifiers for molecules used for training and validation and a configuration file to reproduce training on the same dataset. For the reproduction of Table 1 and Table 2, one can use the tags ‘grappa-benchmark’ and ‘grappa-1.1.0’ respectively. Further details on the package can be found on GitHub.
@@ A.2.1 Graph attentional neural network
Grappa’s graph attentional neural network (Figure A7) is a modification of the transformer architecture [7] for graphs, similar to the Graph Transformer Network [15], but e.g. omits the locality-breaking graph Laplacian eigenvectors as positional encoding. Instead, we encode the graph structure by constraining the attention mechanism to edges of the graph as in GAT [11], which enforces locality: In each attention update only neighboring nodes can influence each other. We use a multi-head dot-product attention mechanism and a 2-layer MLP as feed-forward network, where the hidden feature dimension is four times the node feature dimension.
After initializing the node features $\nu_{i}$ as described below, we apply a single nodewise linear layer followed by an exponential linear unit (ELU) [29],
@@
and biases $b_{1},b_{2}$ are learnable and independent for each layer. We use dropout [30] on the node feature update in (18) and  (19). Finally, we project onto the output dimension $d_{\text{emb}}$ by a linear layer,
@@
As initial node features we choose a one-hot encoding of the atomic number, a one-hot encoding of whether the node is in a loop, a one-hot encoding of the potential loop size and the degree of the node. Since Grappa only predicts bonded parameters, we have to ensure consistency with the nonbonded parameters from the traditional force field, which is why we one-hot encode the traditional force field used for the nonbonded contribution of the respective state if we train on a dataset that contains data from different nonbonded methods. We also include a sinusodial encoding $\tilde{Q}$ of the partial charges $q$ as in the original transformer architecture [7],
$\tilde{Q}_{2i}(q)$ $=\sin\left(\frac{q+2}{4}\cdot 10000^{\frac{-2i}{32}}\right),$ (25)
$\tilde{Q}_{2i+1}(q)$ $=\cos\left(\frac{q+2}{4}\cdot 10000^{\frac{-2i}{32}}\right),$ (26)
with $i\in\{1,\ldots,16\}$, which allows us to describe differently charged conformations of the same molecule without resorting to global or graph-symmetry breaking features like the formal charge.

### v2

@@ 3.2 Grappa is extensible across chemical space
Dataset Test Mols Confs Grappa Espaloma Gaff-2.11 ff99SB-ILDN Mean
Predictor
SPICE-Dipeptide 67 2592 Energy 2.4 3.1 4.5 19.1
Force 5.4 7.8 12.9 38.4
Dipeptides-300K 72 3600 Energy 2.6 4.5 4.1 7.8
Force 5.9 10.9 11.8 43.7
Dipeptides-1000K 72 2160 Energy 5.4 8.6 8.5 18.9
Force 11.6 16.1 17.8 74.3
Non-Capped-Peptides 10 500 Energy 2.2 4.2 4.0 7.2
Force 6.1 12.1 12.6 45.8
Radical-Dipetides 28 272 Energy 3.3 8.7
Force 6.8 41.3
@@
Table 2: Accuracy of Grappa-1.3 with nonbonded parameters from ff99SB, Espaloma 0.3 (Takaba et al., 2023) and established MM force fields on a subset of test molecules from the Grappa-1.3 dataset. The performance on small molecules and RNA is similar to the one reported in Table 1 and can be found in Table A7. As in Table 1, we report the RMSE of centered energies in kcal/mol and the componentwise RMSE of forces in kcal/mol/Å.
If a certain kind of chemistry is not part of the training set of a machine learned force field, predictions can go awry. While the current datasets already cover a significant part of chemical space, extensions should be straightforward and acessible. To demonstrate Grappa’s extensibility, we train it on the Grappa-1.3 dataset, which includes dipeptides sampled at 300 and 1000 K, N- and C-terminal amino acids, dipeptides containing the non-standard residues hydroxyproline and DOPA (dihydroxyphenylalanine), and, demonstrating Grappa’s capability to parametrize rather uncommon molecules, peptide radicals, which play a role in enzyme catalysis (Lebrette et al., 2023) and as mechanoradicals (Zapp et al., 2020).
To create these additional datasets, we calculate DFT energies and gradients of states that were sampled from MD simulations performed with traditional force fields at 300 K and 1000 K, where we increase the temperature to 1000 K in between the samplings of 300 K-states to enrich conformational diversity, as described in A.5.
In Table 2, we report Grappa’s accuracy on the Grappa-1.3 dataset. Again, Grappa outperforms the baselines Amber99SB-ILDN and Gaff-2.11 on all peptide datasets, also for states sampled at 1000 K, indicating that Grappa is robust under conformational perturbations that lead to out-of-equilibrium states. For the radical peptides, there is no baseline since, to the best of our knowledge, Grappa is the first MM force field that covers this part of chemical space. On the other datasets from Espaloma, Grappa remains competitive, as can be seen in Table A7.
The extension to the Grappa-1.3 dataset indicates that Grappa can learn new chemistries accurately without negatively affecting the performance on the original datasets. Due to Grappa’s limited complexity of input features – only connectivity, atomic numbers and partial charges – the chemical space accessible for parameterization is vast. As long as a molecule has a fixed connectivity and the MM energy functional is a reasonable approximation, Grappa can be trained to predict parameters for it.
3.3 Grappa is compatible with multiple nonbonded parameter schemes
Charge Model Energy RMSE Force RMSE
ff99SB 2.55 $\pm$ 0.05 5.85 $\pm$ 0.06
CHARMM36 2.58 $\pm$ 0.07 6.01 $\pm$ 0.07
AM1-BCC 2.79 $\pm$ 0.07 6.13 $\pm$ 0.07
Table 3: Test accuracy of Grappa-1.3 on the Dipeptides-300K dataset in kcal/mol(/Å) with sets of nonbonded parameters obtained from ff99SB, CHARMM36 and AM1-BCC charges with Lennard-Jones parameters from the OpenFF-2.2.0 force field.
Grappa-1.3 is trained to predict bonded parameters that are compatible with partial charges from AM1-BCC, ff99SB and CHARMM36. Table 3 shows that, in combination with all of these charge models, Grappa-1.3 can predict energies and forces of dipeptides to higher accuracy than established force fields (Table 2). Also for partial charges from ff99SB with additional Gaussian noise of scale 0.1 e, Grappa-1.3 is able to predict energies and forces accurately, indicating that the model is robust under small changes in the partial charges and that Grappa can thus be used in combination with charge models that are not present in the training data. This can be partly attributed to the fact that, on the datasets considered, nonbonded interactions contribute only a small fraction to the total interaction. For learning or evaluating the quality of nonbonded parameters by accuracy of energies and forces, we expect datasets that include intermolecular interactions to be more suitable.
3.4 Grappa can be combined with established force fields
Learnable contribution RMSE [kcal/mol(/Å)]
Bond Angle Dihedral Energy Force
2.5 5.9
3.0 6.1
3.9 7.4
3.8 9.2
3.5 11.0
4.1 11.8
Table 4: Test accuracy of Grappa models that predict MM parameters for different interaction types. For the non-learnable types, MM parameters are taken from ff99SB-ILDN. The models are trained on the Dipeptides-300K and -1000K datasets and evaluated on Dipeptides-300K with ff99SB partial charges.
Grappa’s modular formulation allows for learning only a certain type of interaction while keeping the others fixed to those of an established force field. To quantify the improvement of Grappa over the widely used protein force field Amber ff99SB-ILDN (Lindorff-Larsen et al., 2010), we train Grappa models that predict only certain types of MM parameters and evaluate their accuracy in Table 4. While for each contribution, replacing the ff99DB-ILDN parameters with those predicted by Grappa improves the accuracy, the bond and angle contributions have the most potential for improvement.
3.5 Grappa is interpretable
MM parameters
In the classical mechanics potential functions used in MM, a physical meaning can be attributed to parameters. Hence, badly assigned parameters could be noticeable, e.g. if a bond is much shorter or stiffer than bonds between similar atoms or the amplitude of a proper dihedral is untypically high. Apart from exposing the parameters in the simulation files, Grappa also provides figures of the parameter distributions and, if available, comparisons to parameters from an established force field upon parametrization. For the case of the protein ubiquitin considered in section A.6.1, parameter distributions and a comparison to ff99-SBILDN are depicted in Figure A11. While the established protein force field and Grappa-1.3 predict similar bond distances, we observe deviations for other types of parameters.
Latent space
To generalize across the broad range of chemical space covered by the training set, Grappa has to learn highly informative atom embeddings that capture the local environment in the molecular graph. For interpreting this latent representation, we apply a two-dimensional principal component analysis on a set of predicted atom embeddings (3) of test molecules (Figure 4b). It turns out that the two principal components can be related to main group and period of the periodic table of elements, whose structure is learned implicitly by Grappa.
4 Force field validation
@@
In this section, we demonstrate that Grappa is on par with established force fields when it comes to MD simulations of proteins over timescales of hundreds of nanoseconds and that it captures the physics that cause stability of protein folds. All simulations mentioned in this section were performed with GROMACS, for which the Grappa package111 https://github.com/graeter-group/grappa provides a command line interface, briefly described in A.1. Using this interface (or a similar one for OpenMM), large biomolecular systems with over 100,000 atoms are parameterized within minutes on a CPU (Figure A6).
@@ 4.1 Grappa keeps large proteins stable during MD
A key quality check for a protein force field is if a native protein structure remains stable over time during an MD simulation. QM methods and E(3) equivariant neural networks suffer from the problem that while their end-to-end predictions are accurate, they might fail in yielding stable simulations because of diverging gradients of the potential energy surface when extrapolating beyond regions covered by the training dataset. Particularly over long timescales, the system might drift out of the range of validity of the model and small errors might accumulate, leading to instabilities of the system and a potential crash of the simulation. MM with its interpretable, physics-inspired energy function, with parameters which are conformation-independent, ensures a higher degree of stability, and we indeed did not observe simulation crashes.
@@
However, since it is not ensured a priori that macrostates such as protein folds remain stable during MD, we next assess Grappa’s capability of keeping protein folds close to their experimentally determined structure. As example system, we use ubiquitin (PDB ID: 1UBQ), which is a protein with 76 amino acids whose fold contains a beta-sheet and an alpha-helix. We perform a 200 nanoseconds MD simulation of ubiquitin in aqueous solution with Grappa and Amber ff99SB-ILDN (Lindorff-Larsen et al., 2010), for which details are given in A.6.1. We find that the C-alpha RMSD from the initial state is bounded by about 4 Å in 200 ns simulation with both Grappa and Amber ff99SB-ILDN (Figure 5c), indicating that the folded state of the protein is stable. Structural fluctuations on timescales of up to 2 ns are of similar magnitude (Figure 5b).
Figure 5: (a) The protein ubiquitin with color-coded sequence position. (c) The mean C-alpha root mean square deviation (RMSD) and its 25th and 75th percentile of 1000 random pairs of frames that are separated by the time difference $\Delta t$. (b) C-alpha RMSD from the initial state during MD simulation of ubiquitin in water with Grappa.
4.2 Grappa is orders of magnitude more resource-efficient than E3 equivariant models
Efficiency of force fields is crucial; otherwise the computational cost for simulating large systems on long timescales can prohibit their application in simulations of many systems of interest. To demonstrate Grappa’s efficiency, which it inherits from molecular mechanics, and to showcase its capability of jointly parametrizing RNA and proteins, we simulate the virus STMV (The Amber Project, 2023) in solution – a system with approximately one million atoms, visualized in Fig. A9 (a).
A recently proposed machine learned force field relying on an E(3) equivariant neural network with state-of-the-art accuracy and efficiency, Allegro (Musaelian et al., 2022), has been shown to be capable of performing near-quantum-accuracy simulations of systems of unprecedented size. This is achieved by using a strictly local architecture, which allows parallelization across thousands of GPUs. Since they are not constrained by the energy functional of MM, E3 equivariant models like Allegro, NequIP (Batzner et al., 2022) or MACE (Batatia et al., 2023) have greatly improved accuracy but are orders of magnitude more expensive than MM force fields like Grappa.
For the system at hand, Allegro achieves a performance of 106 timesteps per second on 4,000 A100 GPUs (Kozinsky et al., 2023). For Grappa, at reduced accuracy and incapable of describing topological changes directly, we measure a performance of 101 timesteps per second on a single A100 GPU in GROMACS (Abraham et al., 2015) without system-specific optimizations.
@@ 5 Conclusions
With Grappa, we propose a machine learning framework for molecular mechanics force fields with state-of-the-art accuracy that can be used seamlessly in established MD engines. Our experiments show that Grappa is transferable to large biomolecules like proteins and viruses, which sets the stage for large scale simulations at improved accuracy, with the same computational efficiency as established MM force fields.
Unlike most traditional MM force fields, Grappa is not specialized to a certain chemical species, but offers consistent parameterization of molecules with different chemistries, such as a wide range of small molecules, peptides and nucleotides, with a single model at a higher level of accuracy than previous MM force fields. Grappa not only enhances existing parameter sets, but also reaches previously inaccesible regions of chemical space, such as non proteinogenic amino acids or protein radicals.
Since current state-of-the-art E(3) equivariant neural network potentials like Allegro are several orders of magnitude more expensive than molecular mechanics, we regard machine learned molecular mechanics approaches like Grappa as suitable method for simulating large systems, especially if computational ressources are limited or when it is not necessary to simulate all parts of a system at quantum accuracy.
6 Outlook
While Grappa achieves high computational efficiency through molecular mechanics, it also inherits its fundamental limitations, including restricted accuracy and the inability to directly describe chemical reactions. However, Grappa is well suited for efficient reparametrization of large molecules that undergo local topological changes induced by chemical reactions, e.g. in kinetic Monte Carlo simulations (Rennekamp et al., 2020). This is due to both, Grappa’s finite field-of-view, which ensures locality on the molecular graph, as well as the lack of hand-crafted chemical features, which allows to apply the parametrization to cut-out regions of molecules.
Grappa is a general and versatile framework to obtain MM parameters for a broad range of molecules and can be seamlessly extended to new chemistries. To this end, the Grappa package implements workflows for retraining on the provided and custom new datasets.
While we consider the consistency with established nonbonded parameters an advantage of Grappa, an extension of the framework to nonbonded parameters is a straightforward next step. It would render Grappa fully independent from traditional MM force fields and could potentially improve its accuracy further.
@@ 7 Reproducibility statement
Grappa is released as open source software under the GNU General Public License v3.0  [^2] https://github.com/graeter-group/grappa. along with the Grappa dataset and SMILES strings used for the train-val-test partition and configuration files to reproduce Table 1 and Table 2.
Repositories for reproducing the MD simulations 333 https://github.com/LeifSeute/validate-grappa and creating the Grappa-1.3 datasets 444 https://github.com/LeifSeute/grappa-data-creation are publicly available.
@@ Acknowledgements
This research was supported by the Klaus Tschira Foundation. The project has also received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101002812) [F.G.]. We thank Kai Riedmiller and Jannik Buhr for application and software related discussions, Fabian Grünewald, Camilo Aponte-Santamaria and Vsevolod Viliuga for discussions on future applications and Daniel Sucerquia for conversations on the role of internal coordinates in MM.
@@ A.1 The Grappa package
Grappa utilizes PyTorch (Paszke et al., 2019) and DGL (Wang et al., 2020) in its implementation. The package is released as open source software under the GNU General Public License v3.0 and is available on GitHub 555 https://github.com/graeter-group/grappa and on PyPi:
pip install grappa-ff
@@
grappa_gmx -f topol.top -o new_topol.top
@@
# download a pretrained grappa model ff = OpenmmGrappa.from_tag(’grappa-1.3’)
@@
Pretrained released models can be accessed by using a tag; the model weights are downloaded automatically from the respective release on GitHub. Released models also contain a dictionary with results on test datasets, a list of identifiers for molecules used for training and validation and a configuration file to reproduce training on the same dataset. Further details on the package can be found on GitHub.
@@ A.2.1 Graph attentional neural network
Grappa’s graph attentional neural network (Figure A7) is a modification of the transformer architecture (Vaswani et al., 2023) for graphs, similar to the Graph Transformer Network (Yun et al., 2020), but e.g. omits the locality-breaking graph Laplacian eigenvectors as positional encoding. Instead, we encode the graph structure by constraining the attention mechanism to edges of the graph as in GAT (Veličković et al., 2018), which enforces locality: In each attention update only neighboring nodes can influence each other. We use a multi-head dot-product attention mechanism and a 2-layer MLP as feed-forward network, where the hidden feature dimension is four times the node feature dimension.
After initializing the node features $\nu_{i}$ as described below, we apply a single nodewise linear layer followed by an exponential linear unit (ELU) (Clevert et al., 2016),
@@
and biases $b_{1},b_{2}$ are learnable and independent for each layer. We use dropout (Srivastava et al., 2014) on the node feature update in (18) and  (19). Finally, we project onto the output dimension $d_{\text{emb}}$ by a linear layer,
@@
[…]